At locations A and B the electric potential has the values 1.41 V and 5.27 V, respectively. You release a proton from rest at one of these locations and it passes through the other one. At which location should you release the particle? What is its speed when it passes through the other location? Number m/s Repeat the same question, but this time using an electron instead of a proton. O A Number m/s
Answer #1

a) Proton should be released at A point.

from conservation of energy

rac{m*v^2}{2}=e*Delta V

v(2 * 1.6 * 10-19 * (5.27-1.41)/(2.67 * 10-27))

v=2.14*104 m/s

b) Electron should be released at B point

v=sqrt(2*1.6*10^{-19}*(5.27-1.41)/(9.1*10^{-31}))

v=1.16*106